# VQ-SGen: A Vector Quantized Stroke Representation for Creative Sketch Generation

### v2

@@ Abstract
This paper presents VQ-SGen, a novel algorithm for high-quality creative sketch generation. Recent approaches have framed the task as pixel-based generation either as a whole or part-by-part, neglecting the intrinsic and contextual relationships among individual strokes, such as the shape and spatial positioning of both proximal and distant strokes. To overcome these limitations, we propose treating each stroke within a sketch as an entity and introducing a vector-quantized (VQ) stroke representation for fine-grained sketch generation. Our method follows a two-stage framework - in stage one, we decouple each stroke’s shape and location information to ensure the VQ representation prioritizes stroke shape learning. In stage two, we feed the precise and compact representation into an auto-decoding Transformer to incorporate stroke semantics, positions, and shapes into the generation process. By utilizing tokenized stroke representation, our approach generates strokes with high fidelity and facilitates novel applications, such as text or class label conditioned generation and sketch completion. Comprehensive experiments demonstrate our method surpasses existing state-of-the-art techniques on the CreativeSketch dataset, underscoring its effectiveness. The code and model will be made publicly available upon publication.
@@ 1 Introduction
To address these limitations, we propose the first stroke-level sketch generation approach that treats each stroke as an independent entity [34, 22] and encodes it using a novel vector-quantized (VQ) representation. This compact, discrete representation enhances the learning process of the generative model by efficiently capturing essential stroke shape variations while minimizing redundancy. Despite the discrete nature of the representation, we observe semantic-aware clustering within the stroke code space, which lays an ideal foundation for the generator to sample new strokes. Furthermore, our bespoke generator incorporates the shape, semantic, and position into its generation process, effectively samples new sketches in the compressed VQ space, and accurately reasons out the structural and semantic relationship between strokes, resulting in sketches that are both appealing and structurally coherent.
@@
  • To the best of our knowledge, we propose the first stroke-level approach, VQ-SGen, for the challenging creative sketch generation task.
@@ 6 Conclusion
In this paper, we have presented a novel approach for creative sketch generation. Our method can effectively create sketches with appealing appearance and coherent structures. We believe that stroke VQ representations can be used in other sketch-related tasks, and hope to inspire future research in sparse stroke representation learning.
@@ Appendix B Training and Inference Details
Implementation details. We use Adam optimizer [12] for training both the VQ-Representation and Gen-Transformer, while the learning rates are set as $10^{-4}$ for the former and $10^{-5}$ for the latter. We use step decay for VQ-Representation with a step size equal to 10 and do not apply learning rate scheduling for Gen-Transformer. We train our network on 2 NVIDIA V100 GPUs. The VQ-Representation network training takes around 20 hours with a batch size of 64. The Gen-Transformer is trained until convergence taking around 10 hours with a batch size of 8. The maximum stroke length $N$ is 20 for Creative Birds and 35 for Creative Birds to make sure it fits all the sketches. Stroke latent embedding $\bm{e}_{i}^{s}$ is a vector in $\mathbb{R}^{256}$. The number of codes $V$ = 8192, and the code dimensions of both codebooks are $512$. Stroke category number $C=8$ for Creative Birds and 17 for Creative Creatures. The balancing weight $\alpha=0.8$.
@@ Appendix E Further discussions
Limitations. Our method has the following limitations: First, we did not extensively explore all combinations of codebook hyperparameters, so our current configuration may not be globally optimal. This could impact the quality of our model’s generated results. For example, in Fig. A5(a), the wings exhibit stroke disconnection and blurring issues. Efficient hyperparameter search could potentially lead to improvements. Second, due to the decoupling of shape and position, we occasionally observe that the generated bounding box may fail to cover all strokes, resulting in a “hard cropping” of the outermost stroke. The strokes in the wings and head in Fig. A5(b) have experienced varying degrees of “hard cropping”. Introducing a coverage regularizer could help address this issue. We leave the two as future work.
@@ Appendix F Detailed Network Configuration
We also illustrate the detailed network architecture of our tokenizing network, which is used for learning the shape or location codebook. The input is a feature matrix in $\mathbb{R}^{N\times 256}$ composed of shape or location embeddings within a sketch. A Conv1d layer is first employed for feature extraction, followed by a MaxPool layer or ConvTranspose1d layer to compress or expand the feature dimensions. Specifically, the encoder comprises 8 layers grouped into three blocks, each characterized by distinct feature dimensions (i.e., 256, 256, and 512), resulting in a feature of $\mathbb{R}^{N\times 512}$. The “x3” and “x2” of each block in Fig. A7 mean that the block is repeated 3 or 2 times. Subsequently, the feature corresponding to each stroke is replaced with the nearest code and then fed into the decoder. Both decoder branches share the encoder structure. The final output aims to reconstruct the latent input.

### v3

@@ Abstract
This paper presents VQ-SGen, a novel algorithm for high-quality creative sketch generation. Recent approaches have framed the task as pixel-based generation either as a whole or part-by-part, neglecting the intrinsic and contextual relationships among individual strokes, such as the shape and spatial positioning of both proximal and distant strokes. To overcome these limitations, we propose treating each stroke within a sketch as an entity and introducing a vector-quantized (VQ) stroke representation for fine-grained sketch generation. Our method follows a two-stage framework - in stage one, we decouple each stroke’s shape and location information to ensure the VQ representation prioritizes stroke shape learning. In stage two, we feed the precise and compact representation into an auto-decoding Transformer to incorporate stroke semantics, positions, and shapes into the generation process. By utilizing tokenized stroke representation, our approach generates strokes with high fidelity and facilitates novel applications, such as text or class label conditioned generation and sketch completion. Comprehensive experiments demonstrate our method surpasses existing state-of-the-art techniques on the CreativeSketch dataset, underscoring its effectiveness.
@@ 1 Introduction
To address these limitations, we propose a novel stroke-level sketch generation approach that treats each stroke as an independent entity [34, 22] and encodes it using a novel vector-quantized (VQ) representation. This compact, discrete representation enhances the learning process of the generative model by efficiently capturing essential stroke shape variations while minimizing redundancy. Despite the discrete nature of the representation, we observe semantic-aware clustering within the stroke code space, which lays an ideal foundation for the generator to sample new strokes. Furthermore, our bespoke generator incorporates the shape, semantic, and position into its generation process, effectively samples new sketches in the compressed VQ space, and accurately reasons out the structural and semantic relationship between strokes, resulting in sketches that are both appealing and structurally coherent.
@@
  • We propose a novel stroke-level approach, VQ-SGen, for the challenging creative sketch generation task.
@@ 6 Conclusion
In this paper, we have presented a novel approach for creative sketch generation. Our method can effectively create sketches with an appealing appearance and coherent structures. We believe that stroke VQ representations can be used in other sketch-related tasks and hope to inspire future research in sparse stroke representation learning.
Acknowledgment.
The authors would like to thank Oisin Mac Aodha and Ankan Bhunia for proofreading the early draft and the valuable discussions, Haocheng Yuan and Yi Yang for helping prepare comparisons with Diffsketcher. CL was supported by a gift from Adobe.
@@ Appendix B Training and Inference Details
Implementation details. We use Adam optimizer [12] for training both the VQ-Representation and Gen-Transformer, while the learning rates are set as $10^{-4}$ for the former and $10^{-5}$ for the latter. We use step decay for VQ-Representation with a step size equal to 10 and do not apply learning rate scheduling for Gen-Transformer. We train our network on 2 NVIDIA V100 GPUs. The VQ-Representation network training takes around 20 hours with a batch size of 64. The Gen-Transformer is trained until convergence taking around 10 hours with a batch size of 8. The maximum stroke length $N$ is 20 for Creative Birds and 35 for Creative Creatures to make sure it fits all the sketches. Stroke latent embedding $\bm{e}_{i}^{s}$ is a vector in $\mathbb{R}^{256}$. The number of codes $V$ = 8192, and the code dimensions of both codebooks are $512$. Stroke category number $C=8$ for Creative Birds and 17 for Creative Creatures. The balancing weight $\alpha=0.8$.
@@ Appendix C Location Codebook Probing
Figure A4: More DiffSketcher comparison [37] with varying numbers of strokes and CFG weights.
@@ Appendix D Creativity Comparison with Diffsketcher
In order to generate more abstract sketches comparable to ours, two extra hyperparameters can be tuned in DiffSketcher - the number of strokes and the classifier-free guidance (CFG) weight. As shown in Fig. A4, we gradually increase the number of strokes as well as the CFG weight in two examples. When the stroke count is low (10 or 20 strokes), their sketches are too abstract to be recognizable. The smiling bird with 35 strokes starts focusing on realistic head details, but it falls short of creativity.
@@ Appendix E Further discussions
Limitations. Our method has the following limitations: First, we did not extensively explore all combinations of codebook hyperparameters, so our current configuration may not be globally optimal. This could impact the quality of our model’s generated results. For example, in Fig. A6(a), the wings exhibit stroke disconnection and blurring issues. Efficient hyperparameter search could potentially lead to improvements. Second, due to the decoupling of shape and position, we occasionally observe that the generated bounding box may fail to cover all strokes, resulting in a “hard cropping" of the outermost stroke. The strokes in the wings and head in Fig. A6(b) have experienced varying degrees of “hard cropping". Introducing a coverage regularizer could help address this issue. We leave the two as future work.
@@
Point sequence encoding Following ContextSeg [34], we use stroke bitmap sequences rather than point sequences to represent sketches. The latter yields inferior reconstructions (see Fig.5 in [34]). We conducted further statistical analysis in our generation scenario (it is segmentation in ContextSeg). Specifically, we replace the CNN with Sketchformer [24] (a Transformer), and train the VQ-VAE and our generator. In terms of reconstruction, Sketchformer only yields Acc=0.44/Rec=0.32, which is far below ours (Acc=0.96/Rec=0.97). As for the generation, our advantages are more obvious (theirs vs. ours): 75.42 vs. 15.78 (FID↓), 12.31 vs. 18.92 (GD↑), 0.16 vs. 0.53 (CS, close to and $>$0.45 is better), confirming the advantage of bitmap encoding.
@@ Appendix F Detailed Network Configuration
We also illustrate the detailed network architecture of our tokenizing network, which is used for learning the shape or location codebook. The input is a feature matrix in $\mathbb{R}^{N\times 256}$ composed of shape or location embeddings within a sketch. A Conv1d layer is first employed for feature extraction, followed by a MaxPool layer or ConvTranspose1d layer to compress or expand the feature dimensions. Specifically, the encoder comprises 8 layers grouped into three blocks, each characterized by distinct feature dimensions (i.e., 256, 256, and 512), resulting in a feature of $\mathbb{R}^{N\times 512}$. The “x3" and “x2" of each block in Fig. A8 mean that the block is repeated 3 or 2 times. Subsequently, the feature corresponding to each stroke is replaced with the nearest code and then fed into the decoder. Both decoder branches share the encoder structure. The final output aims to reconstruct the latent input.